OpenStudy (anonymous):

write the quadratic equation whose roots are -2 and 4, and whose leading coefficients is 1.

OpenStudy (midhun.madhu1987):

Roots of the quadratic equation are given as : -2 and 4. Sum of roots = -2 + 4 = 2 Product of Roots = -2 * 4 = -8

OpenStudy (midhun.madhu1987):

Now the quadratic equation whose sum and product of roots are given is: